### v1

@@ 2 Related work
Numerous methods exist for selecting the most informative regions from an image, including expectation maximization[34], majority voting [2], wake-sleep algorithm [4], sampling from self-attention or certainty maps[37, 39, 30], and Bayesian optimal experiment design[32]; yet, recently the most predominant solution is using reinforcement learning algorithms, such as variants of the Policy Gradient[48, 27] Deep Q-Learning[6] or Actor-Critic[33].
@@ 3.1 Adaptive glimpse sampling
Let $X$ be an unobserved scene to explore. We assume $X$ to be a rectangle within the Cartesian coordinate system. A glimpse $G$ is a square region within $X$ that can be observed by a camera, specified by the position of its top-left corner $(x,y)$ and its size $d$ (camera field of view). Furthermore, we define glimpse scale as: $z=(d-d_{\text{min}})/(d_{\text{max}}-d_{\text{min}})$, where $d_{\text{max}}$ and $d_{\text{min}}$ are constants denoting the maximum and minimum size of a glimpse that can be captured. Intuitively, a scale of 00 corresponds to the maximum camera zoom level and a scale of $1$ to the widest view possible. Finally, let $C=(x,y,z)$ be the coordinates of glimpse $G$ and constant $d_{\text{cam}}\times d_{\text{cam}}$ denote the sampling resolution of $G$, i.e., the resolution glimpses obtained from the camera sensor.
@@ 5.1 Tasks
Reconstructing the entire scene from observed glimpses verifies comprehensive scene understanding. In Tab. 1 we group the results by pixel percentage (fraction of pixels of the original input image known to the model). Our approach outperforms existing methods by a large margin. In particular, with only 6% of the pixels seen, AdaGlimpse performs better than other methods that had over 18% of the image visible to them. Note that unlike baseline methods, we train our model only on ImageNet-1k without fine-tuning for each evaluated dataset. Qualitative outcomes in Fig. 5 showcase AdaGlimpse producing reconstructions with a higher level of detail, reproducing more objects compared to baseline methods.
@@
Results for multi-class classification on the ImageNet-1k dataset can be found in Tab. 2. AdaGlimpse outperforms all prior methods, achieving $77.54\%$ accuracy compared to $76.13\%$ of the best baseline method STAM [33]. With early exploration termination after reaching $85\%$ probability of predicted class, it requires over $40\%$ less pixels to match STAM accuracy. Visualizations of glimpse selection for classification are presented in Fig. 4. With an early stopping probability threshold of $75\%$, it is able to classify $224\times 224$ images using only a few $32\times 32$ glimpses of $4$ patches each.
@@ Acknowledgments
This paper has been supported by the Horizon Europe Programme (HORIZON-CL4-2022-HUMAN-02) under the project "ELIAS: European Lighthouse of AI for Sustainability", GA no. 101120237. This research was funded by National Science Centre, Poland (grant no. 2023/49/N/ST6/02465). The research was supported by a grant from the Faculty of Mathematics and Computer Science under the Strategic Programme Excellence Initiative at Jagiellonian University. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016558.

### v2

@@ 2 Related work
Numerous methods exist for selecting the most informative regions from an image, including expectation maximization[36, 51], majority voting [2], wake-sleep algorithm [4], sampling from self-attention or certainty maps[39, 41, 32], and Bayesian optimal experiment design[34]; yet, recently the most predominant solution is using reinforcement learning algorithms, such as variants of the Policy Search[50, 29, 28] Deep Q-Learning[6, 7] or Actor-Critic[35].
@@ 3.1 Adaptive glimpse sampling
Let $X$ be an unobserved scene to explore. We assume $X$ to be a rectangle within the Cartesian coordinate system. A glimpse $G$ is a square region within $X$ that can be observed by a camera, specified by the position of its top-left corner $(x,y)$ and its size $d$ (camera field of view). Furthermore, we define glimpse scale as: $z=(d-d_{\text{min}})/(d_{\text{max}}-d_{\text{min}})$, where $d_{\text{max}}$ and $d_{\text{min}}$ are constants denoting the maximum and minimum field of view of an agent camera. Intuitively, a scale of 00 corresponds to the maximum camera zoom level and a scale of $1$ to the widest view possible. Finally, let $C=(x,y,z)$ be the coordinates of glimpse $G$ and constant $d_{\text{cam}}\times d_{\text{cam}}$ denote the sampling resolution of $G$, i.e., the resolution glimpses obtained from the camera sensor.
@@ 5.1 Tasks
Reconstructing the entire scene from observed glimpses verifies comprehensive scene understanding. In Tab. 1 we group the results by pixel percentage (fraction of pixels of the original input image known to the model). Our approach outperforms existing methods by a large margin. In particular, with only 6% of the pixels seen, AdaGlimpse performs better than other methods that had over 18% of the image visible to them. Note that unlike baseline methods, we train our model only on ImageNet-1k without fine-tuning for each evaluated dataset. Qualitative outcomes in Fig. 5 showcase AdaGlimpse producing reconstructions with a higher level of detail, reproducing more objects compared to baseline methods. All models except AttSeg have backbone networks pre-trained on ImageNet-1k; AttSeg’s pre-training details were not disclosed.
@@
Results for multi-class classification on the ImageNet-1k dataset can be found in Tab. 2. AdaGlimpse outperforms all prior methods, achieving $77.54\%$ (±$(\pm 0.18)$ accuracy compared to $76.13\%$ of the best baseline method STAM [35]. With early exploration termination after reaching $85\%$ probability of predicted class, it requires over $40\%$ less pixels to match STAM accuracy. Visualizations of glimpse selection for classification are presented in Fig. 4. With an early stopping probability threshold of $75\%$, it is able to classify $224\times 224$ images using only a few $32\times 32$ glimpses of $4$ patches each.
@@ Acknowledgments
This paper has been supported by the Horizon Europe Programme (HORIZON-CL4-2022-HUMAN-02) under the project "ELIAS: European Lighthouse of AI for Sustainability", GA no. 101120237. This research was funded by National Science Centre, Poland (grant no. 2023/49/N/ST6/02465, 2022/47/B/ST6/03397, 2022/45/B/ST6/02817, and 2023/50/E/ST6/00469). The research was supported by a grant from the Faculty of Mathematics and Computer Science under the Strategic Programme Excellence Initiative at Jagiellonian University. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016558.